Grade Select GradeMagical Mathematics[Interesting Approach]

Find the last number of the following number series
10, 8, 16,13, 39, 35, ___
a.) 75
b.) 100
c.) 130
d.) 140
Explain

Rakhi Agrawal
11 Years agoGrade Select Grade

2 Answers

ng29
11 Years ago
answer is 140
to do such type of questions try to find some pattern like addition, subtraction . multipliy or divide
in this que 10-8=2       8*2=16      16-13=3  
    13*3=39        39-35=4     35*4=140
 
approve if useful
 
grenade
11 Years ago
 
D because of  10-8=2    8*2=16    16-13=3    13*3=39        39-35=4     35*4=140